ImSonique:

What is the value of x in the equation 2x + 3y = 36, when y = 6? 8 9 27 36

Xaophu:

9

Xaophu:

Step 1: Simplify both sides of the equation. 2x+18=36 Step 2: Subtract 18 from both sides. 2x+18−18=36−18 2x=18 Step 3: Divide both sides by 2. so x=9